OpenStudy (anonymous):

What is the discount rate for $50 headphones that are on sale for $42? A. 4.2% B. 8% C. 16% D. 84%

OpenStudy (kayley01):

D because you do 50 times 0.84 if you do that it equals 42!

OpenStudy (anonymous):

ok thank you!

OpenStudy (kayley01):

your welcome!

OpenStudy (anonymous):

Yep D

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

\[\frac{1}{50} (50-42) 100=16 \% \]

OpenStudy (anonymous):

it was not d...

OpenStudy (anonymous):

it was c...

OpenStudy (anonymous):

That's what I thought

OpenStudy (anonymous):

lol

OpenStudy (kayley01):

oh I am sorry I thought it was D

OpenStudy (anonymous):

R u doing a test?

OpenStudy (kayley01):

does it say why it was D?

OpenStudy (anonymous):

cuz u have to do 50x.16=8 and then 50-8=42

OpenStudy (kayley01):

oh ok